I have two questions:
Solve the system and find the point of the intersection.
The sum of two numbers is 2. If one number is subtracted from the other, their difference is 8.

Thanks in Advance

Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)

I will answer one for you.
Reminder :
Three of The site rules :
One problem per submission
No similar problems
No more than 4 submissions daily
You have violated them all.
a+b=2
a-b=8
add the two equations
2a=10
a=5
b=-3

Answer by stanbon(75887)

I have two questions:
Solve the system and find the point of the intersection.
The sum of two numbers is 2. If one number is subtracted from the other, their difference is 8.
---
x + y = 2
x - y = 8
---
Add and solve for "x":
2x = 10
x = 5
---
Since x + y = 2, y = 2-5 = -3
--------------------------------------
The sum of three times a first number and twice a second number is 8. If the second number is subtracted from twice the first number, the result is 3.
3x + 2y = 8
2x - y = 3
------
y = 2x-3
Substitute for "y" and solve for "x":
3x + 2(2x-3) = 8
3x + 4x - 6 = 8
7x = 14
x = 2
---
Since y = 2x-3
y = 2*2-3 = 1
====================
Cheers,
Stan H.